# Christian Peukert [^1] Corresponding author. Florian Abeillon Jérémie Haese Franziska Kaiser Alexander Staub

### v1

@@ 3.2 The Unsplash research program
3.2.1 Two-tiered access to training data for AI
@@
In August 2020, Unsplash released a dataset containing 25,000 “nature-themed” and “featured” photos available for commercial and non-commercial use, including for the training of AI. This dataset is called the LITE dataset. In addition, the platform released a dataset covering the whole collection of images hosted on the platform (hereafter the FULL dataset) made available exclusively for non-commercial research purposes. Access to the LITE dataset happens with the click of a button, access to the FULL dataset is subject to an application process with a thorough manual review process. The datasets do not contain the images per se (they can still be downloaded quite effortlessly through an API), but costly-to-compute metadata such as keyword tags, depicted landmarks, color distributions, etc. By releasing the two flavors of the dataset, Unsplash essentially treated their entire community with their image (metadata) being available for AI research. However, the images that were selected for the LITE dataset receive a much stronger treatment because access is quick, free, and subject to less restrictive usage permissions.
3.2.2 The selection of images in LITE
@@
The LITE dataset was created on June 25, 2020, and, therefore, is a subset of all images available on Unsplash up until that date. The images to be included in the dataset were selected using the following database query: [^6] See https://github.com/unsplash/datasets/issues/55 for details on the sampling procedure as expressed by Timothy Carbone, Head of Data at Unsplash, who runs the research dataset program. In an email conversation with us, he shared the database query he used to select the images in LITE and clarified further: “The Lite dataset’s only objective is to give a preview of the content that’s in the Full dataset. Here are the motivations behind its content: (1) Focusing on a single keyword helps the end-user accessing more depth of content, rather than breadth. This is allegedly more helpful for things like training AI and will hopefully provide a better ‘trial’ experience for the dataset. (2) Making sure we only choose curated content helps with lowering the risk of the content being removed from the platform and makes the dataset easier on the eye.”
1SELECT DISTINCT photos.id FROM photos
2JOIN tags ON tags.photo_id = photos.id
3WHERE tags.keyword = ’nature’ -- The photo is tagged with "nature"
5 AND photos.curated = TRUE -- photo has been curated by the editorial team
6 AND photos.status = 1 -- photo is still available on the platform
7LIMIT 25000
Unsplash uses third-party computer vision algorithms to automatically add tags to images, most prominently Amazon’s Rekognition service. [^7] See https://unsplash.com/blog/the-data-stack-at-unsplash/ and https://docs.aws.amazon.com/rekognition/latest/APIReference/API_DetectLabels.html for details. These auto-generated keywords come with a confidence score which ranges from 0 to 1.
Curated images are photos that were chosen by a human curator at Unsplash to be “special”, i.e., unique perspective, exploration of light and materials, creative use of color, inventive framing, evoking a mood, usefulness for reuse. These photos were either featured on the front page of the website or featured for ten days under a specific topic (e.g.: “nature”). [^8] See https://unsplash.com/blog/how-we-choose-what-photos-to-feature-on-the-unsplash-homepage/.
@@
The fact that there was no specific ordering in the SQL query, i.e. no ORDER BY clause, leads to an unpredictable row order of the results. The row order depends on the database’s internal storage and retrieval mechanisms, which can be influenced by factors like the structure of the database’s indexes and how the query optimizer chooses to execute the query. [^9] The documentation of PostgreSQL, which is the database system Unsplash uses, states: “When using LIMIT, it is a good idea to use an ORDER BY clause that constrains the result rows into a unique order. Otherwise you will get an unpredictable subset of the query’s rows […] You don’t know what ordering unless you specify ORDER BY. […] If ORDER BY is not given, the rows are returned in whatever order the system finds fastest to produce.”, see https://www.postgresql.org/docs/current/sql-select.html and “If sorting is not chosen, the rows will be returned in an unspecified order. The actual order in that case will depend on the scan and join plan types and the order on disk, but it must not be relied on. A particular output ordering can only be guaranteed if the sort step is explicitly chosen.”, see https://www.postgresql.org/docs/current/queries-order.html.
@@
Note: White-robust standard errors in parentheses. $p<0.10$, ∗∗ $p<0.05$∗∗ $p<0.01$
@@
Table 1 shows that the order of images in the LITE dataset cannot be explained by image characteristics such as image age (i.e. time since upload), image popularity (i.e. views and downloads) or whether the user has chosen the keyword “nature” to describe the image. Also, user characteristics such as the total sum of views and downloads of the users’ images (UserPopularity) and the time since the first upload of the user (AccountAge), do not explain the order of images in the LITE dataset. [^10] Note that we do not have user-level information about 14 users, because their accounts were deleted before we can observe them in FULL 1.0.1, the earliest FULL dataset we have access to. These users have 63 images in the LITE dataset. Hence, the number of observations in column 2 is smaller than in column 1. Only the confidence score associated with the automatically assigned keyword “nature” is correlated with the sort order. The higher the confidence score of an image, the higher up it appears on the list. This is consistent with PostgreSQL using a descending order to generate a list of all keywords that meet the confidence condition. For example, conditional on meeting the threshold, keywords with a confidence score of 0.99 would appear before keywords with a confidence score of 0.90. However, it is worth stressing that the regression can only explain little variation in the rank order, as shown by the low value of the adjusted $R^{2}$.
@@
With this information, we can construct a control group of images, for example, consisting of those images that would have shown up below row 25,000 had the query not used a LIMIT condition. Based on the results in Table 1, images would be listed further up, the higher the confidence value of the computer vision algorithm for the keyword “nature”. For technical reasons, however, this does not necessarily mean that those images monotonically convey more of the concept “nature”. They can also convey the same amount of “nature” but less of other concepts. In multi-label classification algorithms using Deep Learning and Convolutional Neural Networks, as deployed in the DetectLabels service of Amazon Rekognition, the calculation of confidence scores is a nuanced process. [^11] The exact technical underpinnings of Rekognition are not publicly disclosed, but a slide deck from a System Dev Engineer at Amazon Webservices gives some indication about the service’s architecture, see https://iptc.org/download/events/phmdc2017/IPTC-PhMdC2017-AHornsby-AmazonRekognition.pdf. The last layer typically applies the softmax function, transforming logits – the network’s final layer output, representing unnormalized log probabilities – into a probability distribution across the predicted labels. Softmax is non-linear implying that changes in logits can be disproportionally reflected in confidence scores. Moreover, the confidence score for a given label is not only a function of its own logit but also depends on the logits of all other labels. Therefore, an increase in the confidence score of “nature”, disproportionally so at the high end of the range, may result from a larger logit of the concept itself or from smaller logits of other concepts.
@@ 3.2.3 Announcement and stakeholders’ awareness of the program
The release of the research datasets was announced publicly on the Unsplash blog on August 6, 2020.121212See https://unsplash.com/blog/the-unsplash-dataset/. In this announcement, Unsplash mentioned the target users for the dataset: “We’re releasing the most complete high-quality open image dataset ever, free for anyone to use to further research in machine learning, image quality, search engines, and more.” They also describe the differences between the two versions: “We’re releasing the data in two versions: a LITE dataset available for commercial and noncommercial usage, and the FULL dataset available for noncommercial usage.” Importantly, they also acknowledge that the dataset is not a one-shot project: “As the Unsplash library continues to double in size every year, we’ll continue updating the dataset with new fields and new images.” Every user who subscribes to the Unsplash newsletter, which is the default option when signing up for an account, receives periodic emails summarizing announcements on the Unsplash blog. Therefore, we have strong reasons to believe that the average user was aware of the research dataset program. Users were not able to opt out, but were, of course, free to delete affected images or their user accounts. Users also did not receive monetary or other compensation for having images included in the LITE nor the FULL dataset. The release of the Unsplash dataset was well received by the AI community. The GitHub repository, which provides documentation and sample code, was starred 2,300 times, and there is an active discussion with the community via issues and pull requests. [^13] See https://github.com/unsplash/datasets. In general, according to PapersWithCode, data from Unsplash has been used in at least 24 computer science papers for problems such as Image Augmentation, Image Generation, Image Inpainting and Outpainting, and Text-to-Image Generation. [^14] See https://paperswithcode.com/datasets?q=unsplash&v=lst&o=match. Altogether, these are strong reasons to believe that the release of the LITE dataset was sufficiently salient to Unsplash users to allow us to detect treatment effects.
4 Data and methods
4.1 Available information
We have access to rich metadata on more than 4.9 million images, which represents all images uploaded to Unsplash since May 2013 and still available in June 2020 and/or in May 2023 (FULL dataset). In particular, we combine information from datasets released in August 2020 (but created on June 25, 2020) and in May 2023.151515To be precise, we have access to datasets in version 1.0.0 (LITE), and versions 1.1.0 and 1.2.1 (FULL). See https://github.com/unsplash/datasets/blob/master/CHANGELOG.md. We observe the user account that has uploaded the image, the upload date, whether an image was curated, a list of keywords associated with each image, and metadata on camera gear for photos. We also observe whether a user has chosen to display a badge on their Unsplash page that says “Available for hire” and for a subset of users (those who chose to reveal this information in their Unsplash profile), we also know their Instagram account name and have collected information on their upload history on Instagram. [^16] This information is not available in the LITE or FULL datasets. We, therefore, do not have time-varying information, but we were able to get a snapshot on February 5, 2024, from which we observe the for-hire status and Instagram information as of that day of about 11,352 out of 12,052 users from our preferred sample of treated and control users. We observe Instagram activity for a subset of 1,084 users. Not all Unsplash users reveal their Instagram accounts on their Unsplash profiles. Further, technical restrictions require us to limit ourselves to the 20 most recent posts of a user. We exclude users for whom we only observe posts in the after period as we do not know whether this is because the account didn’t exist in the before period, whether there were no posts in the before period, or whether the user is so active that we can only see posts from the after period in our snapshot of the 20 most recent posts. As a result, our sample is likely biased towards less active users with relatively old Instagram accounts. Further, we know whether a user has joined the Unsplash+ program as a contributor. [^17] We do not know when a particular user joined the Unsplash+ program. However, we have this information as of April 23, 2024, approximately 2.5 years after the program’s launch.
@@ 4.1 Available information
Using information about the camera gear that the photographer used, we flag users as professionals if at least one of the following conditions is fulfilled for any photo they uploaded before the release of the LITE dataset. First, the exposure time is longer than 5 seconds, which cannot be executed by low-end cameras – including most smartphones – and without a tripod. Second, a photo was shot with a focal length of more than 250, which typically requires expensive professional lenses. Third, a photo is shot on a camera from a mid-to-high-end brand, i.e., not on a smartphone. [^18] We do this based on the following list of brands: Hasselblad, Leica, Rollei, RED, Blackmagic, Voigtlander, Canon, Nikon, Fuji, Sigma, Olympus, Ricoh, Minolta, Konica, Yashica, Kodak, Agfa, Noritsu, Epson, DJI.
4.2 Treatment and control groups
We exploit the fact that Unsplash released a subset of their entire platform contents as a freely available LITE dataset. Because we want to study the behavioral responses of users when their work is made available as a training dataset for AI research, our treatment group is comprised of all 8,298 user accounts that have at least one photo in the LITE dataset. [^19] Users can change account names on Unsplash. Using the fact that the IDs of photos remain unique and cannot be changed, we can track the same user even if they changed account names. There are several ways to construct a control group. First, we could simply take all users who had uploaded at least one image before the LITE dataset was created but did not have any images included in the LITE dataset. Second, we could take the set of users whose images fulfilled the opposite of the selection criteria of the original database query discussed in section 3.2.2. That is, we take users who had not uploaded nature-themed and curated photos at the time when the LITE dataset was created. Third, we could take all users whose images fulfilled the same criteria as in the original database query but were not included in the LITE dataset. This approach comes closest to resembling the perfect counterfactual for users in the LITE dataset, making it our preferred specification. This specification of the control group includes 3,754 users. In the robustness section, we also report results from the two former approaches, showing that the results are consistent, but our preferred approach yields the most conservative estimates.
4.3 Panel construction
@@
To analyze dynamics over time, we need to create a panel structure at the user level. We construct two separate panels. First, while we have information on the upload date of every image, we do not have high-frequency information on whether an image is still available on the platform. We have two snapshots from which we can construct a dataset that tells us whether a user account we observe carrying active images in June 2020 still has any available images in May 2023.
Second, to study contribution dynamics in detail, we aggregate the data at the monthly level and count the number of uploads of every user within that time period. To create a balanced panel, we impute zeroes in months where we do not observe any activity.
4.4 Analysis of variety and novelty
To investigate whether contributions to Unsplash change after the release of the LITE dataset, we compare the flow of new uploads to the stock of images one year before the launch of the research program. In particular, to capture variety as a quality measure of a training dataset, we measure how similar each new upload is relative to all existing images. To capture novelty, we count the number of images in the stock of data that are very similar to a focal new upload.
We then compare how variety and novelty evolve over time for treated users and control users, before and after the release of the LITE dataset. We do so based on the keywords that users assign to their uploaded images. The average image has 10.28 keywords. We use natural language processing techniques to assess the similarity between images based on their associated keywords in 4 steps.
  1. 1.
    Aggregate keywords for each image and train a Word2Vec model on these aggregated keywords to generate a vector representation for each unique keyword. Herein, the Word2Vec model is configured to create 100-dimensional vectors for each keyword, with a context window of 5 words and a minimum word frequency threshold of 1, ensuring that even rare keywords are included in the model.
  2. 2.
    Apply the Word2Vec model to each image, converting the keywords into vectors, and then calculate the average vector to represent the image’s semantic content.
  3. 3.
    Calculate cosine similarity scores between the vector representation of a particular image and the vector representations of all existing images.
  4. 4.
    Compute average similarity across all images and count the number of images that exceed a specified similarity threshold, respectively.
We perform steps 1 and 2 for the stock of images uploaded by users that eventually will be treated, as well as for the stock of images uploaded by users in the control group. We define the stock as all images uploaded before June 25, 2019. We perform steps 1-4 for all images uploaded by treated and control users after June 25, 2019. This allows us to compare new uploads starting one year before the start of Unsplash’s research program to new uploads after the release of the first LITE dataset, across users in the treatment and in the control group. We report 0.8 and 0.9 as the threshold level for very similar images in step 4, where cosine similarity scores are defined on the interval $[-1,1]$, with higher values corresponding to higher similarity between images.
4.5 Identification strategy and econometric model

### v2

@@ 3.1 The Unsplash research program
Unsplash is a stock image platform hosting approximately 6 million photos and illustrations, submitted by roughly 360,000 contributors. According to Similarweb, Unsplash is among the top 5 most visited websites in the “photography” category globally, receiving about 30 million visits each month. Over the past decade, Unsplash has recorded a cumulative total of 1 trillion image views, coming from tens of millions of individual users and thousands of partner websites and applications. Unlike other stock photography websites, the vast majority of images on Unsplash are made available under a permissive license, which allows them to be used freely for both non-commercial and commercial purposes. Compared to other websites that make images available under similar licenses (e.g., Wikimedia), contributions on Unsplash are of comparatively higher quality. Partially, this is enabled by rigorous moderation efforts of an editorial team that hand-selects images that are prominently shown on the website. As of May 2023, the curation team has selected about 300,000 images for curation, which represents about 6% of all images available on the platform.
@@
In August 2020, Unsplash started a research program and released a dataset containing 25,000 images available for commercial and non-commercial use, including for the training of AI. This dataset is called the LITE dataset. In addition, the platform released a dataset covering all images hosted on the platform (hereafter the FULL dataset) made available exclusively for non-commercial research purposes. Access to the LITE dataset happens with the click of a button, while access to the FULL dataset is subject to an application process with a thorough manual review process. [^3] By releasing the two flavors of the dataset, Unsplash essentially treated their entire community with their image (metadata) being available for AI research. However, the images that were selected for the LITE dataset receive a much stronger treatment because access is quick, free, and subject to less restrictive usage permissions. As we show in section 4.1.4, users who did not have images in the LITE dataset (and therefore only in the FULL dataset), do not seem to change the amount of contributions they make to Unsplash compared to their amount of contributions to Instagram. In contrast, we see a significant difference for users with images in the LITE dataset across Unsplash but not on Instagram. This implies that there is no causal change in behavior of being included in the FULL dataset without being part of LITE.
Both released datasets contain costly-to-compute metadata such as keyword tags, depicted landmarks, color distributions, etc., as well as the URLs to download the full-resolution images. The bulk download of the images themselves is explicitly permitted [^4] In response to the question “I would like to download the images using the provided urls. Do I risk getting blocked for making too many requests/downloads”, the Head of Data at Unsplash writes on GitHub: “you shouldn’t be blocked if you use the provided URLs to download the photos. You can try it out and raise an issue if anything goes wrong.” (See https://github.com/unsplash/datasets/issues/37., straightforward and fast (a sample script to download all images in the LITE dataset is provided in the appendix Listing LABEL:lst:image_download). As a result, the LITE release effectively provides easy access to both metadata and image files in a commonly accepted format for image training datasets. [^5] Out of 38,679 datasets currently available on Huggingface, only 10,885 include image files - see https://huggingface.co/datasets?modality=modality:image and https://huggingface.co/datasets?modality=modality:image&format=format:imagefolder
3.1.1 The selection of images in the LITE dataset
@@
The LITE dataset was created on June 25, 2020, and, therefore, is a subset of all images available on Unsplash until that date. The images to be included in the dataset were selected based on four criteria: (1) images are nature-themed, (2) images have been curated by the editorial team, (3) images are currently available on the platform, (4) the total number of images in the dataset is maximum 25,000.666See https://github.com/unsplash/datasets/issues/55 for details on the sampling procedure as expressed by Timothy Carbone, Head of Data at Unsplash, who runs the research dataset program. In an email conversation with us, he shared the database query he used to select the images in LITE (see Listing LABEL:lst:query) and clarified further: “Focusing on a single keyword helps the end-user access more depth of content, rather than breadth. This is allegedly more helpful for things like training AI and will hopefully provide a better ‘trial’ experience for the dataset. Making sure we only choose curated content helps with lowering the risk of the content being removed from the platform and makes the dataset easier on the eye.” To determine whether an image is nature-themed, Unsplash uses computer vision technology that assigns keywords to images, most prominently Amazon’s Rekognition service. [^7] See https://unsplash.com/blog/the-data-stack-at-unsplash/ and https://docs.aws.amazon.com/rekognition/latest/APIReference/API_DetectLabels.html for details. These auto-generated keywords come with a confidence score, which ranges from 0 to 1. Curated images are selected by a human curator at Unsplash to be “special”, meaning they possess a unique perspective, exploration of light and materials, creative use of color, inventive framing, mood evocation, or usefulness for reuse. These images were either featured on the front page of the website or featured for ten days under a specific topic (e.g.: “nature”). [^8] See https://unsplash.com/blog/how-we-choose-what-photos-to-feature-on-the-unsplash-homepage/.
@@
To pull the data for the LITE dataset, Unsplash used a query without specific ordering in the SQL query, i.e., no ORDER BY clause (see Listing LABEL:lst:query). In PostgreSQL, which is the database system they used at the time, this leads to an unpredictable row order of the results. The row order depends on the database’s internal storage and retrieval mechanisms, which can be influenced by factors like the structure of the database’s indexes and how the query optimizer chooses to execute the query. [^9] The documentation of PostgreSQL, which is the database system Unsplash uses, states: “When using LIMIT, it is a good idea to use an ORDER BY clause that constrains the result rows into a unique order. Otherwise, you will get an unpredictable subset of the query’s rows […] You don’t know what ordering unless you specify ORDER BY. […] If ORDER BY is not given, the rows are returned in whatever order the system finds fastest to produce.”, see https://www.postgresql.org/docs/current/sql-select.html and “If sorting is not chosen, the rows will be returned in an unspecified order. The actual order in that case will depend on the scan and join plan types and the order on disk, but it must not be relied on. A particular output ordering can only be guaranteed if the sort step is explicitly chosen.”, see https://www.postgresql.org/docs/current/queries-order.html.
@@
Note: The dependent variable is the rank order of images in the LITE dataset. White-robust standard errors in parentheses. $p<0.10$, ∗∗ $p<0.05$∗∗ $p<0.01$
@@
Indeed, Table 1 shows that the order of images in the LITE dataset cannot be explained by image characteristics such as image popularity (i.e., views and downloads), whether the user has chosen the keyword “nature” to describe the image (UserKeyword) or image age (i.e., time since upload). [^10] Note that we do not have user-level information about 14 users because their accounts were deleted before we can observe them in FULL 1.0.1, the earliest FULL dataset we have access to. These users have 63 images in the LITE dataset. Hence, the number of observations in column 2 is smaller than in column 1. Also, user characteristics such as the total sum of views and downloads of the users’ images (UserPopularity), the time since the first upload of the user (AccountAge), and the total number of uploaded images before treatment (TotalUploads) do not explain the order of images in the LITE dataset. Only the confidence score associated with the automatically assigned keyword “nature” (AutoKeywordScore) is correlated with the sort order. The higher the confidence score of an image, the higher up it appears on the list. This is consistent with PostgreSQL using a descending order to generate a list of all keywords that meet the confidence condition. For example, conditional on meeting the threshold, keywords with a confidence score of 0.99 would appear before keywords with a confidence score of 0.90. However, it is worth stressing that the regression can only explain little variation in the rank order, as shown by the low value of the adjusted $R^{2}$.
@@
With this information, we can construct a control group of images, for example, consisting of those images that would have appeared below row 25,000 had the query not used a LIMIT condition. Based on the results in Table 1, images would be listed further up, the higher the confidence value of the computer vision algorithm for the keyword “nature”. For technical reasons, however, this does not necessarily mean that those images monotonically convey more of the concept “nature”. They can also convey the same amount of “nature” but less of other concepts. [^11] In multi-label classification algorithms using Deep Learning and Convolutional Neural Networks, as deployed in the DetectLabels service of Amazon Rekognition, the calculation of confidence scores is a nuanced process. While the exact technical underpinnings of Rekognition are not publicly disclosed, a slide deck from a System Dev Engineer at Amazon Webservices gives some indication about the service’s architecture (see https://iptc.org/download/events/phmdc2017/IPTC-PhMdC2017-AHornsby-AmazonRekognition.pdf). The last layer typically applies the softmax function, transforming logits – the network’s final layer output, representing unnormalized log probabilities – into a probability distribution across the predicted labels. Softmax is non-linear implying that changes in logits can be disproportionally reflected in confidence scores. Moreover, the confidence score for a given label is not only a function of its own logit but also depends on the logits of all other labels. Therefore, an increase in the confidence score of “nature”, disproportionally so at the high end of the range, may result from a larger logit of the concept itself or from smaller logits of other concepts.
@@ 3.1.2 Announcement and stakeholders’ awareness of the program
The release of both research datasets was announced publicly on the Unsplash blog on August 6, 2020.121212See https://unsplash.com/blog/the-unsplash-dataset/. In this announcement, Unsplash mentioned the target users for the dataset: “We’re releasing the most complete high-quality open image dataset ever, free for anyone to use to further research in machine learning, image quality, search engines, and more.” They also describe the differences between the two versions: “We’re releasing the data in two versions: a LITE dataset available for commercial and noncommercial usage, and the FULL dataset available for noncommercial usage.”. Every user who subscribes to the Unsplash newsletter, which is the default option when signing up for an account, receives periodic emails summarizing announcements on the Unsplash blog. Therefore, we have reasons to believe that the average user was aware of the research dataset program. Users were not able to opt out, but were, of course, free to delete affected images or their user accounts. Users also did not receive monetary or other compensation for having their works included in the LITE or the FULL dataset. The release of the Unsplash dataset was well received by the AI community. The GitHub repository, which provides documentation and sample code, was starred 2,500 times, and there is an active discussion with the community via issues and pull requests. [^13] See https://github.com/unsplash/datasets. In general, according to PapersWithCode, data from Unsplash has been used in at least 24 computer science papers for problems such as image augmentation, image generation, image inpainting and outpainting, and text-to-image generation. [^14] See https://paperswithcode.com/datasets?q=unsplash&v=lst&o=match. Altogether, these are strong reasons to believe that the release of the LITE dataset was sufficiently salient to Unsplash users to allow us to detect treatment effects. Further, in section 4.2, we provide anecdotal evidence on the stated reactions of the photography community.
3.2 Data and methods
3.2.1 Available information
@@
We have access to rich metadata on more than 4.9 million images, encompassing all images uploaded to Unsplash since May 2013 that remained available in either June 2020 or May 2023 (the FULL dataset). Specifically, we combine information from datasets released in August 2020 (but created on June 25, 2020) and May 2023.151515To be precise, we have access to datasets in version 1.0.0 (LITE), and versions 1.1.0 and 1.2.1 (FULL). See https://github.com/unsplash/datasets/blob/master/CHANGELOG.md. We observe the user account that has uploaded the image, the upload date, whether an image was curated, and a list of keywords associated with each image (both user-specified and generated by computer vision algorithms). Using information about the camera gear that the photographer used, we flag users as professionals if at least one of the following conditions is fulfilled for any image they uploaded before the release of the LITE dataset. First, the exposure time is longer than 5 seconds, which cannot be executed by low-end cameras – including most smartphones – and without a tripod. Second, a photo was shot with a focal length of more than 250, which typically requires expensive professional lenses. Third, a photo is shot on a camera from a mid-to-high-end brand, i.e., not on a smartphone. [^16] We do this based on the following list of brands: Hasselblad, Leica, Rollei, RED, Blackmagic, Voigtlander, Canon, Nikon, Fuji, Sigma, Olympus, Ricoh, Minolta, Konica, Yashica, Kodak, Agfa, Noritsu, Epson, DJI. Another measure of whether a user is likely a professional photographer is whether they have chosen to display a badge on their Unsplash page that says “Available for hire”. Finally, we know whether a user has joined the Unsplash+ program as a contributor (see section 5.1 for more details on the program). [^17] We do not know when a particular user joined the Unsplash+ program. However, we have this information as of April 23, 2024, approximately 2.5 years after the program’s launch.
3.2.2 Treatment and control groups
@@
To analyze user-level contribution dynamics over time, we construct two separate panel datasets. First, although we lack high-frequency information on whether an image is still available on the platform, we have information on the upload date of every image from two snapshots that let us determine whether a user account with active images in June 2020 still has any images available in May 2023. Second, to study contribution patterns in detail, we aggregate data at the monthly level and count the number of uploads of every user within that time period. To create a balanced panel, we impute zeroes in months where we do not observe any activity.
To define treatment and control groups, we exploit the fact that Unsplash released a subset of their entire platform contents as a freely available dataset for commercial use. Because we want to study the behavioral responses of users when their work is made available as a training dataset for commercial AI, our treatment group is comprised of all 8,298 user accounts that have at least one image in the LITE dataset. [^18] Users can change account names on Unsplash. Using the fact that the IDs of images remain unique and cannot be changed, we can track the same user even if they changed account names. The treatment concerns users with images present in the first release of the LITE dataset. Further, while de-facto randomization happened at the image level, mechanically, some users may have a higher chance of having images included in the LITE dataset because the distribution of contributions to the platform is not uniform. We correct for potential implications by estimating models with user-fixed effects that remove user-specific time-invariant variation (such as generally being more active on the platform), as well as by estimating models that weigh users inversely proportional to their pre-existing stock of contributions. There are several ways to construct a control group. First, we could take all users who had uploaded at least one image before the LITE dataset was created but did not have any images included in the LITE dataset. Second, we could take the set of users whose images fulfilled the opposite of the selection criteria discussed in section 3.1.1, i.e., users who had not uploaded nature-themed and curated images when the LITE dataset was created. Third, we could take all users whose images fulfilled the same criteria as in the original database query but were not included in the LITE dataset. This approach comes closest to resembling the perfect counterfactual for users in the LITE dataset, making it our preferred specification. This specification of the control group includes 3,754 users. In the robustness section, we also report results from the two former approaches, showing that the results are consistent, but our preferred approach yields the most conservative estimates.
Further, we can exploit the fact that some users contribute to Unsplash as well as Instagram, an unrelated image-sharing platform that is not affected by Unsplash’s research program. Uploads to Instagram serve as a counterfactual of what upload behavior on Unsplash would have been had they not released the LITE and FULL datasets. As we report in the robustness section, we show that only users whose data was released for commercial usage (i.e. were part of LITE) showed a causal reaction.
3.2.3 Identification strategy and econometric model
Our identification strategy is based on the notion that the images in the LITE dataset are randomly drawn from a superset of all images on the platform that meet the eligibility criteria.